# colorlinks = true, linkcolor = blue, citecolor = blue, urlcolor = blue

### v1

@@ Abstract
Object rearrangement is a significant task for collaborative robots, where they are directed to manipulate objects into a specified goal state. Determining the placement of objects is a major challenge that influences the efficiency of the rearrangement process. Most current methods heavily rely on pre-collected datasets to train the model for predicting the goal position and are restricted to specific instructions, which limits their broader applicability and effectiveness. In this paper, we propose a framework of language-conditioned object rearrangement based on the Large Language Model (LLM). Particularly, our approach mimics human reasoning by using past successful experiences as a reference to infer the desired goal position. Based on LLM’s strong natural language comprehension and inference ability, our method can generalise to handle various everyday objects and free-form language instructions in a zero-shot manner. Experimental results demonstrate that our methods can effectively execute the robotic rearrangement tasks, even those involving long sequential orders.
@@ I INTRODUCTION
Robots have the potential to assist humans with various daily tasks, such as placing objects in desired locations, tidying up toys by gathering them into a designated container and performing household cleaning. These tasks can be categorised as object rearrangement, in which robots manipulate objects to achieve a specified goal state within a physical environment [1]. Effective object rearrangement involves several sub-tasks, including the recognition of object states within the environment, the inference of difference between the current and goal state, and the manipulation of objects accordingly. While there are many advanced solutions for object recognition and manipulation in computer vision, tactile sensing and robotic fields [2][3], we argue that there is a lack of intelligent methods that can perform human-like thinking to infer optimal object placement in rearrangement tasks. Specifically, achieving human’s ability to adapt to changing circumstances and guide flexible behaviour [4][5] by making use of similar past experiences on robots, still needs to be researched.
@@
Currently, most existing methods attempt to learn a mapping from the current position to the goal position [6][7]. Firstly, a large dataset describing the goal position for the rearrangement needs to be collected. Then, the robot can be trained using the current position as the input to predict the goal position with an algorithm or a deep learning model. However, these methods rely heavily on the dataset, making them learn only specific patterns and human instructions from the dataset. When faced with a new environment containing different objects not included in the dataset or new instructions, these methods do not generalise effectively. Most recently, generative models like DALL-E, which is pre-trained on web-scale data, have been used to generate the goal poses in rearrangement to improve the generalisation ability [8]. However, these models can sometimes generate hallucinations, producing random objects and positions that require filtering in a further step.
@@
Figure 1: Learn form the past. In our framework, the robot retrieves past experiences to find the most similar arrangement based on human instructions. By referencing previously successful arrangements, the robot can mimic human-like reasoning with LLMs, allowing it to infer the goal position for rearrangement more effectively.
@@
When people rearrange items according to new instructions, it is natural for them to draw on their past successful experiences [9]. For example, “placing an apple on a plate” may remind them of a similar prior arrangement, such as “putting an orange in a bowl”. In this case, the experience of placing an orange in a bowl can serve as a template, providing scaffolding for the required robot trajectories and positions of both the orange and the bowl. This past experience then guides them in completing the new instructions. As a result, the efficiency of rearrangement tasks is significantly improved, as people distill and transfer knowledge from their previous experiences to new tasks.
@@
Most recently, Large Language Models (LLMs) such as ChatGPT have significantly enhanced multiple fields within robotics, including code generation, task grounding, and human-robot interaction, thanks to their robust natural language understanding and zero/few-shot learning capabilities [10][11]. In this paper, we introduce a framework that links past experiences with new instructions by leveraging natural language comprehension (See Fig. 1). We then make use of an LLM to draw from similar past experiences and perform spatial reasoning to find the position for object placement in rearrangement tasks. As a result, our proposed method allows the robot to mimic human-like reasoning in object rearrangement, and our results demonstrate that the robots can execute rearrangement tasks efficiently, even in tasks with long sequential orders.
@@
    We propose a novel framework based on LLMs that leverages past successful experiences to address the problem of language-conditioned object rearrangement.
@@
    Our method is not limited by the dataset and can be generalised to free-form natural language in a zero-shot manner.
@@
    By emulating human reasoning to learn from past experiences, our approach significantly enhances the performance of rearrangement tasks.
Figure 2: Illustration of the proposed framework. The robot uses SAM for visual perception and CLIP for semantic understandingidentify where and what objects are in the environment. The LLM then associate the most similar past experience with instruction and uses this similar experience as a reference. Finally, a prompt is created with spatial and semantic information, allowing the LLM to predict the goal position for rearrangement.
@@ II Related works
In this section, we will first review works on predicting the goal positions in object rearrangement, followed by discussions of LLM’s applications in robotics. Finally, we will review the Visual Language Model (VLM) which connects visual perception and language instruction in robotics.
@@ II-A Prediction of Goal Position in Rearrangement
The prediction of the goal position (That is the positions of each physical object in the environment) plays an important role in the object rearrangement task. The goal position can be represented in various ways such as through logical predicates, visual images or natural language instructions [1]. In [6][7], the spatial relations of the logical predicates, such as inside, on and next to are modelled between different objects to predict the goal position. Alternatively, for visual image representation, objects are manually arranged into goal states according to human preferences. In this situation, robots can then learn the latent space of user preferences from the image representation and manipulate the objects accordingly [12][13][14].
Additionally, some prior works treat position prediction as a searching task by selecting from a set of sampled poses using graphs. For example, in [15], a GNN is trained to understand the rules of the task, with its nodes representing entities relevant to the task such as objects and goals, which are used to select an appropriate position. In [16], an embodied agent is developed to detect out-of-place objects and infer plausible receptacles through a neural graph memory. However, the methods mentioned above typically require a large dataset that is manually collected for learning purposes, which can be both costly and time-consuming.
Compared to visual goal specification, language offers an intuitive way to describe the goal state. In [17][18], researchers detected objects within the environment and established spatial as well as semantic relationships between them. As a result, robots are able to relocate a specified object with respect to a designated anchor object towards fulfilling a set of given instructions. However, these methods concentrate on learning visual-semantic relationships and are unable to infer the goal position in a manner similar to human reasoning. In our framework, we leverage past successful experiences, similar to how humans do, and use the LLM to perform the spatial reasoning for the language-conditioned rearrangement, for the first of its kind.
@@ II-B LLM and VLM in Language-conditioned Manipulation
Recently, LLMs have become an active area of research in the field of robotics due to their extensive internalised knowledge and chain of thought generation [11]. For example, in [19][16][20], language models are applied to decompose instructions into sequences of sub-steps. Based on the observed environment, the robot can execute the feasible sub-tasks to successfully achieve the instructions. Contrary to decomposing instructions, in [21], the LLM is applied to summarise the rules of arrangement from a small number of examples by leveraging its few-shot learning capabilities. In [10], the LLM is used to generate the planning strategy to enhance the stability of the system. In our framework, we employ the LLM to connect past experiences with new instructions and leverage it for spatial reasoning in rearranging objects, for the first time.
Most lately, Vision-Language Models (VLMs) have become a successful paradigm for aligning visual information with language, enabling robots to perceive the world in a multimodal manner. In [22], an end-to-end model based on VLMs is proposed. This model learns from large amounts of online data to create action commands that demonstrate robust generalisation capabilities in unseen environments. In [23], observations like images and state estimates are integrated into the language embedding space, facilitating more efficient inferences, particularly for sequential decision-making. However, these methods usually use large robotic data for training, which is a bottleneck because of limited data availability. To address this issue, [24] used pre-trained LLM and VLM to infer the affordances and constraints within the environment based on the language instructions provided. This approach enabled the execution of physical robot actions without relying on robotic data. In our framework, we apply the pretrained VLM to align the visual information with semantic information in a similar way and use the LLM together with past experience as a reference to perform spatial reasoning to estimate the goal position in a zero-shot manner.
@@ III Methodology
The main objective of our framework is to leverage knowledge from previous arrangement experience and infer optimal positions for object placement conditioned on new instructions. The pipeline is shown in Fig. 2. First, the robots perform the vision-language grounding by using the semantic understanding provided by the Contrastive Language-Image Pre-Training Model (CLIP) and the vision precision from Segment Anything Model (SAM). This allows the robot to identify where and what the objects are in the environment. Then, we apply the LLM to associate the most relevant experience with the given instruction and use the spatial and semantic information from this similar experience as a reference. Next, we generate a prompt combining the spatial and semantic information of both the observed environment and the referenced experience. Based on this prompt, the LLM can effectively predict the position for object placement in rearrangement tasks, conditioned on the language instruction.
@@ III-A Vision-language Grounding
For language-conditioned rearrangement, vision-language grounding, which is the aligning of each object with the corresponding semantic information, plays a fundamental role in the task. To predict the goal state of each object, we need to create an object-level representation, using the RGB observation of the environment from the camera. Specifically, we employ a pre-trained SAM for segmenting objects in RGB images. The model processes the image and point-based prompts as inputs and accordingly produces the object mask. By sampling single-point prompts across a grid on the image, SAM is able to generate a mask and bounding box for each point.
After generating the masks across a grid, semantic filtering is performed on the image to obtain the desired objects in the language instruction. In order to understand the objects to be manipulated, we first feed the LLM with language instruction and the prompt: “List the objects that are directly involved in the interaction described in the instruction”. In response, the LLM will identify and output the objects related to the instruction. For instance, if the instruction given is “put the apple next to the banana”, the LLM will output the objects involved as [“apple”, “banana”, “others”] where the “others” is given for the irrelevant objects. To align segmented objects with semantic information, the open-vocabulary image classifier CLIP is then used to match the masked images with corresponding names. Finally, an object representation is generated, which includes spatial information such as centre points, masks, bounding box, rotation, and the corresponding object names. Algorithm 3 demonstrates the core idea of the vision-language grounding with the observed RGB images.
@@ III-B Spatial reasoning with retrieval augmentation generation
The key to object rearrangement is identifying the position where the target object should be placed. Due to the strong capabilities of language comprehension and internalised knowledge, LLM has great potential to become a general problem solver, extending beyond writing stories and essays. Based on the spatial information from observed images, such as the objects’ centre points, masks, bounding box, rotation, and semantic details, the LLM can be used to perform spatial reasoning to infer the objects’ target position for the rearrangement task.
@@
However, the dataset used to train the LLM is limited, as many LLMs are trained on corpora from the internet, books, and videos, but lack spatial information. Consequently, LLMs often struggle with spatial reasoning abilities and can not understand the shape accurately [25]. Inspired by the Retrieval Augmentation Generation (RAG) [26], we design a RAG system that allows the LLM to retrieve outer knowledge of past successful arrangement experiences and enhance its generated results. Specifically, this outer knowledge of successful rearrangement includes spatial and semantic information, along with the corresponding instructions. The system identifies the most similar arrangement from the outer knowledge as a reference. As shown in Fig. 2, the LLM scores the similarity between the new instruction and each instruction in past successful experiences. Detailed spatial and semantic information from the most closely matched experience is incorporated into the prompt to facilitate spatial reasoning. An example of a simplified prompt template is shown in Fig. 3. As a result, the LLM can predict the goal position more accurately during spatial reasoning.
{algorithm}
@@
[…]
object_names
[…]
@@ III-C Execute Pick and Place
Once the LLM identifies the object to be moved and the desired position for placement in the pixel coordinate, the robot can pick and place the object for the rearrangement. An RGB-D camera is mounted on the wrist of the robotic arm. Given the object’s initial and target locations in pixel coordinates, we can map these to their real-world coordinate using the depth information provided by the camera. We then employ inverse kinematics and a motion planner to find the path for the manipulation and move the object to complete the rearrangement task.
Figure 4: Examples of successful rearrangements. Four examples of successful rearrangements arranged by humans, with corresponding instructions.
Figure 5: Demonstration of results. The first row shows results from random object placement. The second displays objects arranged in a horizontal line with a certain gap. The third shows results from the proposed method without reference, and the fourth presents results using our proposed method.

### v2

@@ Abstract
Object manipulation for rearrangement into a specific goal state is a significant task for collaborative robots. Accurately determining object placement is a key challenge, as misalignment can increase task complexity and the risk of collisions, affecting the efficiency of the rearrangement process. Most current methods heavily rely on pre-collected datasets to train the model for predicting the goal position. As a result, these methods are restricted to specific instructions, which limits their broader applicability and generalisation. In this paper, we propose a framework of flexible language-conditioned object rearrangement based on the Large Language Model (LLM). Our approach mimics human reasoning by making use of successful past experiences as a reference to infer the best strategies to achieve a current desired goal position. Based on LLM’s strong natural language comprehension and inference ability, our method generalises to handle various everyday objects and free-form language instructions in a zero-shot manner. Experimental results demonstrate that our methods can effectively execute the robotic rearrangement tasks, even those involving long sequences of orders.
@@ I INTRODUCTION
Robots have the potential to assist humans with various daily tasks, such as placing objects in desired locations, tidying up toys by gathering them into a designated container and performing household cleaning. These tasks can be categorised as object rearrangement, in which robots manipulate objects to achieve a specified goal state within a physical environment [1]. Effective object rearrangement involves several sub-tasks, including the recognition of object states within the environment, the inference of difference between the current and goal state, and the manipulation of objects accordingly. While there are many advanced solutions for object recognition and manipulation in computer vision, tactile sensing and robotic fields [2][3], we argue that there is a lack of intelligent and flexible methods that can perform human-like thinking to infer optimal object placement in rearrangement tasks while generalising to unforeseen scenarios not included in the training dataset. This is especially true for intra-class tasks. Currently, most existing methods attempt to learn a mapping from the current position to the goal position [4][5].
@@
In these methods, a large dataset describing the goal position for the rearrangement is collected and used to train deep learning algorithms running on the robot. Using the current position as input, the algorithm predicts the goal position and directs the robot’s movement accordingly. However, these methods rely heavily on the training dataset and makes them learn only specific patterns with specific objects and human instructions from the dataset. This works for inter-class tasks but when faced with a new environment containing different objects not included in the dataset or new instructions (intra-class), these methods do not generalise effectively [6]. Most recently, generative models like DALL-E, which is pre-trained on web-scale data, have been used to generate the goal poses in rearrangement to improve the generalisation ability [7]. However, these models can sometimes generate hallucinations, producing random objects and positions that require filtering in a further step.
@@
Figure 1: Learn from the past. In our framework, the robot retrieves past experiences to find the most similar arrangement based on human instructions. By referencing previously successful arrangements, the robot can mimic human-like reasoning (See Fig. 2), allowing it to infer the goal position for rearrangement more effectively.
Figure 2: Learn from the past in humans. A diagram showing how humans make use of past experiences to guide successful current and future task completions in life long learning [8]. Mental models are used as templates or references for current tasks. Successful mental models are stored for future reuse.
@@
As a result, achieving human’s ability to adapt to changing circumstances and guide flexible behaviour [9][10] through the use of similar past experiences (See Fig. 2), still needs to be researched. When humans respond to new instructions to manipulate or rearrange items both in inter-task and intra-task scenarios, it is natural for them to draw on their past successful experiences [11][8]. For example, the instruction “place an apple on a plate” may remind them of a similar prior arrangement and instruction, such as “put an orange in a bowl”. In this case, the experience of placing an orange in a bowl can serve as a mental model template that provides scaffolding for reasoning about the current task as well as generating the required trajectories and positions for completing the task involving both the apple and the plate. By leveraging past experiences, the completion efficiency of tasks is improved, as people distil and transfer knowledge from their previous experiences to new tasks (See Fig. 2).
@@
Following this inspiration, we introduce a framework that links past experiences with new instructions by leveraging natural language comprehension and reasoning (See Fig. 1). Specifically, we make use of Large Language Models (LLMs) to draw on similar past experiences towards constructing templates for performing spatial reasoning in current rearrangement tasks. As a result, our proposed method allows the robot to mimic human-like flexibility in language conditioned (i.e., human language instructed) object rearrangement tasks while executing the tasks efficiently and accurately, even in tasks with long sequential orders.
@@
    We propose a novel framework that leverages past successful experiences to address the problem of flexible object rearrangement tasks.
@@
    Our framework is not limited by a training dataset and can be generalised to various tasks using free-form natural language instructions in a zero-shot manner.
@@
    By emulating human reasoning in leveraging past experiences, our framework significantly enhances the performance of current rearrangement tasks.
Figure 3: Illustration of the proposed framework. The robot uses SAM for visual perception and CLIP for semantic understanding to identify where and what objects are in the environment. The LLM then associate the most similar past experience with instruction and uses this similar experience as a template and reference. Finally, a prompt is created with spatial and semantic information, allowing the LLM to reason and predict the goal position for rearrangement.
@@ II Related works
In this section, we will first review works on predicting the goal positions in object rearrangement, followed by discussions on LLM’s applications in robotics. Finally, we will review the Visual Language Model (VLM) which connects visual perception and language instruction in robotics.
@@ II-A Prediction of Goal Position in Rearrangement
The prediction of the goal position (that is the positions of each physical object in the environment) plays an important role in the object rearrangement task. The goal position can be represented in various ways such as through logical predicates, visual images or natural language instructions [1]. In [4][5], the spatial relations of the logical predicates, such as inside, on and next to are modelled between different objects to predict the goal position. Additionally, some prior works treat position prediction as a searching task by selecting from a set of sampled poses using graphs. For example, in [12], a GNN is trained to understand the rules of the task, with its nodes representing entities relevant to the task such as objects and goals, which are used to select an appropriate position. In [13], an embodied agent is developed to detect out-of-place objects and infer plausible receptacles through a neural graph memory. However, the methods mentioned above typically require a large dataset that is manually collected for learning purposes, which can be both costly and time-consuming. The approaches above can be combined with visual image detection and representation in which objects are manually arranged into goal states according to human preferences. In these situations, robots learn the latent space of user preferences from the image and manipulate the objects accordingly [14][15][16][17].
Nevertheless, compared to only visual goal specification, language enriches human robot collaboration by offering a more intuitive way to describe the goal state. For example, in [18][19], researchers used vision to detect objects within the environment and established spatial as well as semantic relationships between them. Then robots were able to relocate a specified object with respect to a designated anchor object towards fulfilling a set of given natural language instructions. However, these methods concentrate on learning visual-semantic relationships and are unable to infer the goal position in a manner similar to human reasoning. In our framework, we use LLMs to draw on similar past successful experiences and to perform spatial reasoning in language-conditioned flexible rearrangement of objects. By providing sufficient information regarding the current task and similar previous tasks, including the object’s name, position, etc., the robot is able to construct the appropriate template (or mental model in humans) for comprehension, reasoning and completing the current task. According to our current knowledge, this is the first of its kind.
@@ II-B LLM and VLM in Language-conditioned Manipulation
Recently, LLMs have become an active area of research in the field of robotics due to their extensive internalised knowledge and chain of thought generation [20]. For example, in [21][13][22], language models are applied to decompose instructions into sequences of sub-steps. Based on the observed environment, the robot can execute the feasible sub-tasks to successfully complete the instructions. Contrary to decomposing instructions, in [23], a LLM is applied to summarise the rules of arrangement from a small number of examples by leveraging its few-shot learning capabilities. Additionally, in [24], the proposed method refines long-horizon behaviours by chaining basic actions together with guidance from an LLM. In [25], a wake-sleep framework is applied, where the wake phase employs an LLM-based actor-critic to interact with the environment based on human demonstrations and hints, and the sleep phase clusters experiences with an LLM abstractor. However, these methods focus on reordering basic actions to form new behaviors rather than learning entirely novel skills. In our framework, we employ a LLM to connect past arrangements with new instructions and use it as a reference (or mental model) for spatial reasoning in estimating a more accurate target state, for the first time.
Most lately, Vision-Language Models (VLMs) have become a successful paradigm for aligning visual information with language, enabling robots to perceive the world in a multimodal manner. In [26], an end-to-end model based on VLMs is proposed. This model learns from large amounts of online data to create action commands that demonstrate robust generalisation capabilities in unseen environments. In [27], observations like images and state estimates are integrated into the language embedding space, facilitating more efficient inferences, particularly for sequential decision-making. However, these methods usually use large robotic data for training, which is a bottleneck because of limited data availability. To address this issue, [28] used pre-trained LLM and VLM to infer the affordances and constraints within the environment based on the language instructions provided. This approach enabled the execution of physical robot actions without relying on robotic data. In our framework, we apply the pretrained VLM to align the visual information with semantic information in a similar way and use the LLM together with past experience as a reference to perform spatial reasoning in estimating the goal position in a zero-shot manner.
@@ III Methodology
The main objective of our framework (See Fig. 3) is to leverage knowledge from previous arrangement experience and infer optimal positions for object placement conditioned on new instructions. First, the robot performs vision-language grounding by using semantic understanding provided by the Contrastive Language-Image Pre-Training Model (CLIP). This is supported by vision precision from a Segment Anything Model (SAM). It allows the robot to identify what objects and where they are in the environment. Then, we apply a Retrieval-Augmented Generation (RAG) [29] system to enable the robot to leverage past successful experiences for a rearrangement task. Particularly, we apply the LLM to associate the most relevant experience with the given instruction and use the spatial and semantic information from this similar experience as a reference. Next, we generate a prompt combining the spatial and semantic information of both the observed environment and the referenced experience. Based on this prompt, the LLM can accurately predict the position for object placement in rearrangement tasks, conditioned on the language instruction.
@@ III-A Vision-language Grounding
For language-conditioned rearrangement, vision-language grounding, which is the aligning of each object with the corresponding semantic information, plays a fundamental role in the task. To predict the goal state of each object, we need to create an object-level representation, using the RGB observation of the environment from the camera. Towards this, we apply a pre-trained SAM for segmenting objects in RGB images. The model uses images and point-based prompts as inputs and then produces the segmented object. By sampling single-point prompts across a grid on the image, SAM is able to generate a mask and minimum bounding box for each sampled location. Concretely, given an input RGB image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$, SAM segments object regions $\mathbf{M}_{i}$, each associated with a minimum bounding box $\mathbf{B}_{i}=\left(x_{i},y_{i},w_{i},h_{i},{\theta}_{i}\right)$, where $\left(x_{i},y_{i}\right)$ denotes the centre point of the box, $\left(w_{i},h_{i}\right)$ are the width and height of the box, and ${\theta}_{i}$ is the rotation of the rectangle. By using the SAM, we obtain the masked images and corresponding minimum bounding boxes:
$\mathbf{M}_{i},\mathbf{B}_{i}=\operatorname{SAM}\left(\mathbf{I},\mathbf{p}_{i}\right)$ (1)
where $p_{i}$ represents sampled points on the image in a grid.
After generating the masks across a grid, semantic filtering is performed on the image to obtain the desired objects in the language instruction. In order to understand the objects to be manipulated, we first feed the LLM with language instruction and the prompt $\mathcal{Q}_{obj}$: “List the objects that are directly involved in the interaction described in the instruction”. In response, the LLM identifies and outputs the objects related to the instruction. This is represented as $\mathcal{C}_{{obj}}=\operatorname{LLM}(\mathcal{L},\mathcal{Q}_{obj})$ where $\mathcal{L}$ represents the given instruction. For instance, if the given instruction is “put the apple next to the banana”, the LLM will output the objects involved as [“apple”, “banana”, “others”] where the “others” is given for the irrelevant objects. To align segmented objects with semantic information, the open-vocabulary image classifier CLIP is used to match the masked images with corresponding names. Particularly, the visual features and semantic features are extracted by the visual encoder and the semantic encoders of CLIP respectively. We then compute the cosine similarity between the visual and text (i.e., object name) features:
$s_{i,c}=\frac{\mathbf{v}_{i}\cdot\mathbf{t}_{c}}{\left\|\mathbf{v}_{i}\right\|\left\|\mathbf{t}_{c}\right\|}$ (2)
where $v_{i}$ and $t_{c}$ denote the visual features of segmented objects and semantic features respectively. Finally, we assign the object to the category with the highest similarity, which is denoted as $c_{i}=\arg\max_{c\in\mathcal{C}_{\text{obj }}}s_{i,c}$. We can easily use the corresponding minimum bounding boxes and masked objects to obtain the centroid $m_{i}$ and rotation angle $r_{i}$ of each object. As a result, an object representation $\mathcal{O}=\left\{\left(c_{i},m_{i},r_{i},\mathbf{B}_{i}\right)\right\}_{i=1}^{N}$ is generated, which includes spatial information such as centroid, bounding box, rotation, and the corresponding object names.
@@ III-B Spatial reasoning with retrieval augmentation generation
The key to object rearrangement is identifying the position where the target object should be placed. Based on the spatial information from observed images, the LLM can be used to perform spatial reasoning to infer the objects’ target position for the rearrangement task.
@@
However, the dataset used to train the LLM is limited, as many LLMs are trained on corpora from the internet, books, and videos, but lack spatial information. Consequently, LLMs often struggle with spatial reasoning abilities and can not understand the shape accurately [30].
@@
$\mathbf{P}_{t}=\operatorname{LLM}\left(\mathcal{O},\mathcal{L},\mathcal{E}^{*}\right).$ (3)
Here, $\mathbf{P}_{t}=\left(x_{t},y_{t},r_{t}\right)$ represents the predicted coordinates and the rotation angle for object placement. Specifically, the LLM is designed to maintain the initial rotation angle to simplify the rearrangement, modifying it if a collision or overlap is detected.
@@ III-C Execute Pick and Place
Once the LLM identifies the object to be moved and the desired position $\mathbf{P}_{t}=\left(x_{t},y_{t},r_{t}\right)$ for placement in the pixel coordinate, the robot can pick and place the object for the rearrangement. An RGB-D camera is mounted on the wrist of the robotic arm. Given the object’s initial and target locations in pixel coordinates, we can map these to their real-world coordinates using the depth information provided by the camera. The grasping pose is determined based on the object’s minimum bounding box, where the gripper is aligned perpendicular to the longest edge of the object and centred to the object’s centroid. We then employ inverse kinematics and a motion planner to find the path for the manipulation and move the object to complete the rearrangement task.
Figure 5: Examples of successful rearrangements by humans. Four examples of successful rearrangements arranged by humans, with corresponding instructions.
Figure 6: Demonstration of results. The first row shows results from random object placement. The second displays objects arranged in a horizontal line with a certain gap. The third shows results from the application of Dream2Real [31] while the fourth row is our proposed method without a reference or template (having a similar structure with TidyBot [23]). The fifth row presents results using our complete proposed method with reference.